# Recursively-Constrained Partially Observable Markov Decision Processes

### v2

@@ Abstract
In many problems, it is desirable to optimize an objective function while imposing constraints on some other objectives. A Constrained Partially Observable Markov Decision Process (C-POMDP) allows modeling of such problems under transition uncertainty and partial observability. Typically, the constraints in C-POMDPs enforce a threshold on expected cumulative costs starting from an initial state distribution. In this work, we first show that optimal C-POMDP policies may violate Bellman’s principle of optimality and thus may exhibit unintuitive behaviors, which can be undesirable for some (e.g., safety critical) applications. Additionally, online re-planning with C-POMDPs is often ineffective due to the inconsistency resulting from the violation of Bellman’s principle of optimality. To address these drawbacks, we introduce a new formulation: the Recursively-Constrained POMDP (RC-POMDP), that imposes additional history-dependent cost constraints on the C-POMDP. We show that, unlike C-POMDPs, RC-POMDPs always have deterministic optimal policies, and that optimal policies obey Bellman’s principle of optimality. We also present a point-based dynamic programming algorithm that synthesizes admissible near-optimal policies for RC-POMDPs. Evaluations on a set of benchmark problems demonstrate the efficacy of our algorithm and show that policies for RC-POMDPs produce more desirable behaviors than policies for C-POMDPs.
@@ 1 Introduction
Partially Observable Markov Decision Processes (POMDPs) are powerful models for sequential decision making due to their ability to account for transition uncertainty and partial observability. Their applications range from autonomous driving (Pendleton et al. 2017) and robotics to geology (Lauri, Hsu, and Pajarinen 2023; Wang et al. 2022), asset maintenance (Papakonstantinou and Shinozuka 2014), and human-computer interaction (Chen et al. 2020). Constrained POMDPs (C-POMDPs) are extensions of POMDPs that impose a bound on expected cumulative costs while seeking policies that maximize expected total reward. C-POMDPs address the need to consider multiple objectives in applications such as autonomous rover that may have a navigation task as well as an energy usage budget, or human-AI dialogue systems with constraints on the length of dialogues. However, we observe that optimal policies computed for C-POMDPs exhibit pathological behavior in some problems, which can be opposed to the C-POMDP’s intended purpose.
@@
Consider a rover agent in a cave with two tunnels, A and B, which may have rocky terrains. To model wheel damage, a cost of $10$ is given for traversing through rocky terrain, and 00 otherwise. The agent has noisy observations (correct with a probability of $0.8$) of a tunnel’s terrain type, and hence, has to maintain a belief (probability distribution) over the terrain type in each tunnel. The task is to navigate to the end of a tunnel while ensuring that the expected total cost is below a threshold of $5$. The agent has the initial belief of $0.5$ probability of rocks and $0.5$ probability of no rocks in tunnel $A$, and 00 probability of rocks and $1.0$ probability of no rocks in tunnel $B$.
@@
In this paper, we first provide the key insight that the pathological behavior is caused by the violation of Bellman’s principle of optimality (BPO) of C-POMDP optimal policies. To mitigate the pathological behavior and preserve BPO, we propose an extension of C-POMDPs through the addition of history-dependent cost constraints at each reachable belief, which we call Recursively-Constrained POMDPs (RC-POMDPs). We prove that deterministic policies are sufficient for optimality in RC-POMDPs and that RC-POMDPs satisfy BPO. These results suggest that RC-POMDPs are highly amenable to standard dynamic programming techniques, which is not true for C-POMDPs. Then, we present a simple point-based dynamic programming algorithm to solve RC-POMDPs with scalar costs. Experimental evaluation shows that the pathological behavior is a prevalent phenomenon in C-POMDP policies, and that our algorithm for RC-POMDPs computes polices which obtain expected cumulative rewards competitive with C-POMDPs without exhibiting such behaviors. RC-POMDPs provide a good balance between the BPO-violating expectation constraints of C-POMDPs and constraints on the worst-case outcome, which are overly conservative for POMDPs with inherent state uncertainty.
In summary, this paper contributes (i) a proof that BPO does not always hold in C-POMDPs, (ii) the introduction of RC-POMDPs, a novel extension of C-POMDPs through the addition of history dependent cost constraints, (iii) proofs that all RC-POMDPs satisfy BPO, and have at least one deterministic optimal policy, (iv) a dynamic programming algorithm to solve RC-POMDPs, and (v) a series of illustrative benchmarks to demonstrate the advantages of RC-POMDPs.
@@
Several solution approaches exist for C-POMDPs with expectation constraints (de Nijs et al. 2021). These include offline (Isom, Meyn, and Braatz 2008; Kim et al. 2011; Poupart et al. 2015; Walraven and Spaan 2018) and online methods (Lee et al. 2018; Jamgochian, Corso, and Kochenderfer 2023). These works suffer from the pathological behavior discussed above. Our paper shows that this behavior is rooted in the violation of BPO by C-POMDPs and proposes a new problem formulation that obeys BPO.
BPO violation is also present in fully-observable Constrained MDPs (C-MDPs) with state-action frequency and long-run average cost constraints (Haviv 1996; Chong, Miller, and Adaska 2012). To overcome it, Haviv (1996) proposes an MDP formulation with sample path constraints. In C-POMDPs with expected cumulative costs, this BPO-violation problem remains unexplored. Additionally, adoption of the MDP solution of worst-case sample path constraints would be overly conservative for POMDPs, since they are inherently characterized by state uncertainty. This paper fills that gap by studying the BPO of C-POMDPs and addressing it by imposing recursive cost constraints.
From the algorithmic perspective, the closest work to ours is the C-POMDP point-based value iteration (CPBVI) algorithm (Kim et al. 2011). Similar to this work, Piunovskiy and Mao (2000) use the notion of admissible costs. Samples of admissible costs are used with belief points as a heuristic to improve computational tractability of point-based value iteration for C-POMDPs. However, since CPBVI is designed for C-POMDPs, the synthesized policies by CPBVI may still exhibit pathological behavior. Our notion of history-dependent admissible cost bounds is used as problem constraints, eliminating the pathological behavior.
Our work is related to the concept of recursive feasibility in constrained optimal and model predictive control (Mayne 2014; Löfberg 2012). Recursive constraints are also studied in reinforcement learning to reduce learning instability (Lee, Sedwards, and Czarnecki 2022). The recursive constraints in our new formulation can be viewed as a mechanism to enforce closed-loop recursive feasibility, which is not guaranteed under the typical C-POMDP formulation.
@@ 2 Constrained POMDPs
A Partially Observable Markov Decision Process (POMDP) is a tuple $\mathcal{P}=(S,A,O,T,R,Z,\gamma,b_{0})$, where: $S,A,$ and $O$ are finite sets of states, actions and observations, respectively, $T:S\times A\times S\rightarrow[0,1]$ is the transition probability function, $R:S\times A\rightarrow\mathbb{R}$ is the immediate reward function, $Z:S\times A\times O\rightarrow[0,1]$ is the probabilistic observation function, $\gamma\in[0,1)$ is the discount factor, and $b_{0}\in\Delta(S)$ is an initial belief, where $\Delta(S)$ is the probability simplex (the set of all probability distributions) over $S$.
@@
The evolution of an agent according to a POMDP model is as follows. At each time step $t\in\mathbb{N}_{0}$, the agent has a belief $b_{t}$ of its state $s_{t}$ as a probability distribution over $S$ and takes action $a_{t}\in A$. Its state evolves from $s_{t}\in S$ to $s_{t+1}\in S$ according to probability $T(s_{t},a_{t},s_{t+1})$, and it receives an immediate reward $R(s_{t},a_{t})$ and observation $o_{t}\in O$ according to observation probability $Z(s_{t+1},a_{t},o_{t})$. The agent then updates its belief using Bayes theorem; that is for $s_{t+1}=s^{\prime}$,
$b_{t+1}(s^{\prime})\propto Z(s^{\prime},a_{t},o_{t})\sum_{s\in S}b_{t}(s)T(s,a_{t},s^{\prime}).$ (1)
Then, the process repeats. The agent chooses actions according to a policy $\pi:\Delta(S)\to\Delta(A)$, which maps a belief $b$ to a probability distribution over actions. $\pi$ is called deterministic if $\pi(b)$ is a unitary distribution for every $b\in\Delta(S)$. A policy is typically evaluated according to the expected rewards it accumulates over time. Let $R(b,a)=\mathbb{E}_{s\sim b}[R(s,a)]$ be the expected reward for the belief-action pair $(b,a)$. The expected discounted sum of rewards that the agent receives under policy $\pi$ starting from belief $b_{t}$ is
@@
A Constrained POMDP (C-POMDP) is a tuple $\mathcal{M}=(\mathcal{P},C,\hat{c})$, where $\mathcal{P}$ is a POMDP as in Def. 1, $C:S\times A\rightarrow\mathbb{R}^{n}_{\geq 0}$ is a cost function that maps each state action pair to an $n$-dimensional vector of non-negative costs, and $\hat{c}\in\mathbb{R}^{n}_{\geq 0}$ is an $n$-dimensional vector of expected cost thresholds.
@@
In C-POMDPs, the constraint $V^{\pi}_{C}(b_{0})\leq\hat{c}$ is imposed on the POMDP optimization problem as formalized below.
@@
Unlike POMDPs that have at least one deterministic optimal policy (Sondik 1978), optimal policies of C-POMDPs may require randomization, and hence there may not exist an optimal deterministic policy (Kim et al. 2011).
Next, we discuss why the solutions to Problem 1 may not be desirable and a reformulation of it is necessary.
2.1 Violation of Bellman’s Principle of Optimality
Here, we show that Bellman’s Principle of Optimality (BPO) extended to constrained problems does not generally hold for C-POMDPs. Therefore, optimal policies may not exhibit the optimal substructure property, which makes the employment of standard dynamic programming techniques for Problem 1 difficult [^1] There are approaches that use dynamic progamming-like algorithms for C-POMDPs, such as (Isom, Meyn, and Braatz 2008) and (Kim et al. 2011), but they do not find optimal policies.. A similar observation is discussed in (Haviv 1996) for C-MDPs with expected state-action frequency constraints. We show that this is also the case for C-POMDPs with expected cumulative cost constraints.
BPO (Bellman 1957) states that an optimal policy has the property that whatever the initial state and initial decision are, the remaining decisions must constitute an optimal policy with regard to the state resulting from the first decision.
If BPO holds, then there is an optimal substructure property such that the optimal policy $\pi^{*}$ at $b_{0}$ can be computed by finding the optimal policy for each future (successor) belief state according to the same optimization criterion.
To assess the validity of BPO for C-POMDPs, we first formulate it mathematically. Via a trivial extension from Haviv (1996), the Bellman backup operator $B$ for a C-POMDP is:
$B[V^{\pi}](b_{t})=\max_{a\in A}\Big{[}$ $Q_{R}^{\pi}(b_{t},a)\mid Q^{\pi}_{C}(b_{t},a)\leq\hat{c}\Big{]},\,\forall b_{t}.$ (7)
Let $\textsc{Reach}^{\pi}(b_{0})$ be the set of belief states reachable from $b_{0}$ under policy $\pi$. Then, to obey BPO, $\pi^{*}$ must satisfy:
$B[V^{\pi^{*}}\!](b_{t})=V^{\pi^{*}}\!(b_{t})\qquad\forall b_{t}\in\textsc{Reach}^{\pi^{*}}\!(b_{0}).$ (8)
However, this property does not always hold for C-POMDPs as the following theorem states.
Theorem 1.
Eq. (8) does not always hold for C-POMDPs.
Proof.
We provide a counterexample that violates Eq. (8). Consider the POMDP (depicted as a belief MDP) in Figure 1, which is a simplified version of Example 1. W.l.o.g., let $\gamma=1$. The agent starts at $b_{0}$ with a cost constraint $\hat{c}=5$. Actions $a_{A}$ and $a_{B}$ represent going through tunnels A and B, respectively, and $r$ and $nr$ are the observations that tunnel A is rocky and not rocky, respectively.
@@ 2.1 Violation of Bellman’s Principle of Optimality
Figure 1: Counter-example POMDP (depicted as belief MDP) with the associated reward and cost functions. The action at $b_{3}$ has a reward and cost of 00.
@@
By examining the reward function, we see that action $a_{A}$ returns the highest reward at every belief except $b_{0}$. Action $a_{B}$ returns a higher reward at $b_{0}$. Let $\pi_{A}$ be the policy that chooses $a_{A}$ at every belief, and $\pi_{B}$ the one that chooses $a_{B}$ at $b_{0}$. The cost-values for these policies are $V^{\pi_{A}}_{C}(b_{0})=V^{\pi_{B}}_{C}(b_{0})=5\leq\hat{c}$, and the reward-values are $V^{\pi_{A}}_{R}(b_{0})=12$, $V^{\pi_{B}}_{R}(b_{0})=10$. Note that both policies satisfy the cost constraint and any policy that chooses $a_{B}$ at $b_{1}$ or $b_{2}$, or that randomizes between $\pi_{A}$ and $\pi_{B}$ has a value less than $V^{\pi_{A}}_{R}(b_{0})$; hence, $\pi_{A}$ is the optimal policy. However, when using Eq. (7) to compute the optimal policy at $b_{1}$, the optimal action is $a_{B}$ since $a_{A}$ violates the constraint, i.e., $Q_{C}(b_{1},a_{A})=8>\hat{c}$. This is a clear violation of Eq. (8). ∎
The above theorem shows that C-POMDPs do not necessarily obey BPO. Intuitively, this is because, in C-POMDPs, the optimal decision at a belief state $b_{t}$ needs to be informed by (i) the expected cost incurred up to $b_{t}$, as well as (ii) the expected cost incurred by the other beliefs that share an ancestor node with $b_{t}$ in the belief MDP tree.
2.2 Pathological Behavior of Optimal Policies: Stochastic Self-Destruction

### v3

@@ Abstract
Many sequential decision problems involve optimizing one objective function while imposing constraints on other objectives. Constrained Partially Observable Markov Decision Processes (C-POMDP) model this case with transition uncertainty and partial observability. In this work, we first show that C-POMDPs violate the optimal substructure property over successive decision steps and thus may exhibit behaviors that are undesirable for some (e.g., safety critical) applications. Additionally, online re-planning in C-POMDPs is often ineffective due to the inconsistency resulting from this violation. To address these drawbacks, we introduce the Recursively-Constrained POMDP (RC-POMDP), which imposes additional history-dependent cost constraints on the C-POMDP. We show that, unlike C-POMDPs, RC-POMDPs always have deterministic optimal policies and that optimal policies obey Bellman’s principle of optimality. We also present a point-based dynamic programming algorithm for RC-POMDPs. Evaluations on benchmark problems demonstrate the efficacy of our algorithm and show that policies for RC-POMDPs produce more desirable behaviors than policies for C-POMDPs.
@@ 1 Introduction
Partially Observable Markov Decision Processes (POMDPs) are powerful models for sequential decision making due to their ability to account for transition uncertainty and partial observability. Their applications range from autonomous driving Pendleton et al. [2017] and robotics to geology Lauri et al. [2023], Wang et al. [2022], asset maintenance Papakonstantinou and Shinozuka [2014], and human-computer interaction Chen et al. [2020]. Constrained POMDPs (C-POMDPs) are extensions of POMDPs that impose a bound on expected cumulative costs while seeking policies that maximize expected total reward. C-POMDPs address the need to consider multiple objectives in applications such as autonomous rover that may have a navigation task as well as an energy usage budget, or human-AI dialogue systems with constraints on the length of dialogues. However, we observe that optimal policies computed for C-POMDPs exhibit pathological behavior in some problems, which can be opposed to the C-POMDP’s intended purpose.
@@
Consider a rover agent in a cave with two tunnels, A and B, which may have rocky terrains. Traversing tunnel A has a higher expected reward than traversing tunnel B. To model wheel damage, a cost of $10$ is given for traversing through rocky terrain, and 00 otherwise. The agent has noisy observations (correct with a probability of $0.8$) of a tunnel’s terrain type, and hence, has to maintain a belief (probability distribution) over the terrain type in each tunnel. The task is to navigate to the end of a tunnel while ensuring that the expected total cost is below a threshold of $5$. The agent has the initial belief of $0.5$ probability of rocks and $0.5$ probability of no rocks in tunnel $A$, and 00 probability of rocks and $1.0$ probability of no rocks in tunnel $B$.
@@
In this paper, we first provide the key insight that the pathological behavior is caused by the violation of the optimal substructure property over successive decision steps, and hence violation of the standard form of Bellman’s Principle of Optimality (BPO). To mitigate the pathological behavior and preserve the optimal substructure property, we propose an extension of C-POMDPs through the addition of history-dependent cost constraints at each reachable belief, which we call Recursively-Constrained POMDPs (RC-POMDPs). We prove that deterministic policies are sufficient for optimality in RC-POMDPs and that RC-POMDPs satisfy BPO. These results suggest that RC-POMDPs are highly amenable to standard dynamic programming techniques, which is not true for C-POMDPs. RC-POMDPs provide a good balance between the BPO-violating expectation constraints of C-POMDPs and constraints on the worst-case outcome, which are overly conservative for POMDPs with inherent state uncertainty. Then, we present a point-based dynamic programming algorithm to approximately solve RC-POMDPs. Experimental evaluation shows that the pathological behavior is a prevalent phenomenon in C-POMDP policies, and that our algorithm for RC-POMDPs computes polices which obtain expected cumulative rewards competitive with C-POMDPs without exhibiting such behaviors.
In summary, this paper contributes (i) an analysis that C-POMDPs do not exhibit the optimal substructure property over successive decision steps and its consequences, (ii) the introduction of RC-POMDPs, a novel extension of C-POMDPs through the addition of history-dependent cost constraints, (iii) proofs that all RC-POMDPs have at least one deterministic optimal policy, satisfy BPO, and the Bellman operator has a unique fixed point under suitable initializations, (iv) a dynamic programming algorithm for RC-POMDPs, and (v) a series of illustrative benchmarks to demonstrate the advantages of RC-POMDPs.
@@
Several solution approaches exist for C-POMDPs with expectation constraints de Nijs et al. [2021]. These include offline Isom et al. [2008], Kim et al. [2011], Poupart et al. [2015], Walraven and Spaan [2018], Kalagarla et al. [2022], Wray and Czuprynski [2022] and online methods Lee et al. [2018], Jamgochian et al. [2023]. These works suffer from the unintuitive behavior discussed above. This paper shows that this behavior is rooted in the violation of optimal substructure by C-POMDPs and proposes a new problem formulation that obeys BPO.
BPO violation has also been discussed in fully-observable Constrained MDPs (C-MDPs) with state-action frequency and long-run average cost constraints Haviv [1996], Chong et al. [2012]. To overcome it, Haviv [1996] proposes an MDP formulation with sample path constraints. In C-POMDPs with expected cumulative costs, this BPO-violation problem remains unexplored. Additionally, adoption of the MDP solution of worst-case sample path constraints would be overly conservative for POMDPs, which are inherently characterized by state uncertainty. This paper fills that gap by studying the BPO of C-POMDPs and addressing it by imposing recursive expected cost constraints.
From the algorithmic perspective, the closest work to ours is the C-POMDP point-based value iteration (CPBVI) algorithm Kim et al. [2011]. Samples of admissible costs, defined by Piunovskiy and Mao [2000] for C-MDPs, are used with belief points as a heuristic to improve computational tractability of point-based value iteration for C-POMDPs. However, since CPBVI is designed for C-POMDPs, the synthesized policies by CPBVI may still exhibit pathological behavior. In this paper, we formalize the use of history-dependent expected cost constraints and provide a thorough analysis of it. We show that this problem formulation eliminates the pathological behavior of C-POMDPs.
@@ 2 Constrained POMDPs
A Partially Observable Markov Decision Process (POMDP) is a tuple $\mathcal{P}=(S,A,O,T,R,Z,\gamma,b_{0})$, where: $S,A,$ and $O$ are finite sets of states, actions and observations, respectively, $T:S\times A\times S\rightarrow[0,1]$ is the transition probability function, $R:S\times A\rightarrow[R_{min},R_{max}]$, for $R_{min},R_{max}\in\mathbb{R}$, is the immediate reward function, $Z:S\times A\times O\rightarrow[0,1]$ is the probabilistic observation function, $\gamma\in[0,1)$ is the discount factor, and $b_{0}\in\Delta(S)$ is an initial belief, where $\Delta(S)$ is the probability simplex (the set of all probability distributions) over $S$.
@@
The evolution of an agent according to a POMDP model is as follows. At each $t\in\mathbb{N}_{0}$, the agent has a belief $b_{t}$ of its state $s_{t}$ as a probability distribution over $S$ and takes action $a_{t}\in A$. Its state evolves from $s_{t}\in S$ to $s_{t+1}\in S$ according to $T(s_{t},a_{t},s_{t+1})$, and it receives an immediate reward $R(s_{t},a_{t})$ and observation $o_{t}\in O$ according to observation probability $Z(s_{t+1},a_{t},o_{t})$. The agent then updates its belief using Bayes theorem; that is for $s_{t+1}=s^{\prime}$,
$b_{t+1}(s^{\prime})\propto Z(s^{\prime},a_{t},o_{t})\sum_{s\in S}T(s,a_{t},s^{\prime})b_{t}(s).$ (1)
Then, the process repeats. Let $h_{t}=\{a_{0},o_{0},\cdots,a_{t-1},o_{t-1}\}$ denote the history of the actions and observations up to but not including time step $t$; thus, $h_{0}=\emptyset$. The belief at time step $t$ is therefore $b_{t}=P(s_{t}\mid b_{0},h_{t})$. For readability, we do not explicitly include $b_{0}$, as all variables are conditioned on $b_{0}$.
The agent chooses actions according to a policy $\pi:\Delta(S)\to\Delta(A)$, which maps a belief $b$ to a probability distribution over actions. $\pi$ is called deterministic if $\pi(b)$ is a unitary distribution for every $b\in\Delta(S)$. A policy is typically evaluated according to the expected rewards it accumulates over time. Let $R(b,a)=\mathbb{E}_{s\sim b}[R(s,a)]$ be the expected reward for the belief-action pair $(b,a)$. The expected discounted sum of rewards that the agent receives under policy $\pi$ starting from belief $b_{t}$ is
@@
A Constrained POMDP (C-POMDP) is a tuple $\mathcal{M}=(\mathcal{P},C,\hat{c})$, where $\mathcal{P}$ is a POMDP as in Def. 1, $C:S\times A\rightarrow\mathbb{R}^{n}_{\geq 0}$ is a cost function that maps each state action pair to an $n$-dimensional vector of non-negative costs, and $\hat{c}\in\mathbb{R}^{n}_{\geq 0}$ is an $n$-dimensional vector of expected cost thresholds from the initial belief state $b_{0}$.
@@
In C-POMDPs, the constraint $V^{\pi}_{C}(b_{0})\leq\hat{c}$, where $\leq$ refers to the component-wise inequality, is imposed on the POMDP optimization problem as formalized below.
@@
Unlike POMDPs that have at least one deterministic optimal policy Sondik [1978], optimal policies of C-POMDPs may require randomization, and hence there may not exist an optimal deterministic policy Kim et al. [2011].
Next, we discuss why the solutions to Problem 1 may not be desirable and an alternate formulation is necessary.
2.1 Optimal Substructure Property
A problem has the optimal substructure property if an optimal solution to the problem contains optimal solutions to its subproblems Cormen et al. [2009]. Additionally, Cormen et al. note that these subproblems must be independent of each other. If this holds for Problem 1, then the optimal policy $\pi^{*}(b_{0})$ at $b_{0}$ can be computed recursively by finding the optimal policy $\pi^{*}(h_{t})$ for each successive history for the same planning problem. Thus, a natural subproblem to Eq. (6) is the history-based subproblem $(\mathcal{M},h_{t})$, with $\pi^{*}(h_{t})=\arg\max_{\pi}V_{R}^{\pi}(h_{t})\text{ s.t. }V_{C}^{\pi}\left(b_{0}\right)\leq\hat{c}$ [^1] Constraining $V_{C}^{\pi}(h_{t})\leq\hat{c}$ also violates the property as the constraint is defined only at $b_{0}$.. We show that this subproblem violates the optimal substructure property, which makes the employment of standard dynamic programming techniques difficult [^2] Some approaches use dynamic programming (Isom et al. [2008], Kim et al. [2011]), but they do not find optimal policies..
Since the constraint of Eq. (6) is defined only at $b_{0}$, the subproblem at $h_{t}$ must consider the expected cumulative cost of the policy from $b_{0}$. It is not enough to compute the expected total cost obtained from $b_{0}$ to $h_{t}$, as an optimal cost-value from $h_{t}$ depends on cost-values of other subproblems. We illustrate this with an example. Consider the POMDP (depicted as a belief MDP) in Figure 1, which is a simplified version of Example 1. W.l.o.g., let $\gamma=1$. The agent starts at $b_{0}$ with constraint $\hat{c}=5$. Actions $a_{A}$ and $a_{B}$ represent going through tunnels A and B, and $r$ and $nr$ are the observations that tunnel A is rocky and not rocky, respectively.
@@ 2.1 Optimal Substructure Property
Figure 1: Counter-example POMDP with associated reward and cost functions. The action at $b_{3}$ has 00 reward and cost.
@@
By examining the reward function, we see that action $a_{A}$ returns the highest reward everywhere except $b_{0}$. Action $a_{B}$ returns a higher reward at $b_{0}$. Let $\pi_{A}$ be the policy that chooses $a_{A}$ at every belief, and $\pi_{B}$ the one that chooses $a_{B}$ at $b_{0}$. The cost-values for these policies are $V^{\pi_{A}}_{C}(b_{0})=V^{\pi_{B}}_{C}(b_{0})=5\leq\hat{c}$, and the reward-values are $V^{\pi_{A}}_{R}(b_{0})=12$, $V^{\pi_{B}}_{R}(b_{0})=10$. Note that both policies satisfy the constraint and any policy that chooses $a_{B}$ at $b_{1}$ or $b_{2}$, or that randomizes between $\pi_{A}$ and $\pi_{B}$ has value less than $V^{\pi_{A}}_{R}(b_{0})$; hence, $\pi_{A}$ is the optimal policy. However, when planning at $b_{1}$, i.e., $h_{1}$, it is impossible to decide that $a_{A}$ is optimal without first knowing that action $a_{A}$ at $h_{2}$ incurs $2$ cost and is optimal. The decisions at $b_{1}$ and $b_{2}$ cannot be computed separately as subproblems.
To get around this dependence, we can include information about how much cost the policy incurs at other subproblems and how much cost policies can incur from $h_{t}$, obtaining a policy-dependent subproblem $(\mathcal{M},h_{t},\pi)$. This subproblem definition exhibits the optimal substructure property only if we relax the restriction of subproblems being independent. Nonetheless, the optimal solution to a subproblem $(\mathcal{M},h_{t},\pi)$ is only guaranteed to be optimal for the full problem if an optimal policy $\pi^{*}$ is already provided.
2.1.1 Pathological Behavior : Stochastic Self-Destruction
A main consequence of history-dependent subproblems violating the optimal substructure property and instead requiring policy-dependent subproblems is that optimal policies may exhibit unintuitive behaviors during execution.